OpenStudy (anonymous):

evaluate d followin base3 numbers.....212-121+222

OpenStudy (anonymous):

212 with base 3 is 2*3^2+1*3^1+2*3^0=2*9+3+2=23 and so on

Directrix (directrix):

121 = 1 * 3^2 + 2 * 3^1 + 1 * 3^0 = 9 + 6 + 1 = 16.

Directrix (directrix):

222 = 2* 3^2 + 2*3^1 + 2^3^0 = 2*9 + 2*3 + 2*1 = 26

OpenStudy (anonymous):

@Directrix please leave something for Trustted to do :-)

Directrix (directrix):

I did leave something for him to do but didn't wait long enough for him to respond. Point taken.

OpenStudy (anonymous):

yes, because you are helping, not doing the work instead of a person.